For the following decision table, the highest value for the equally likely criterion is _______ ;this occurs with alternative _______ .


States of Nature
Alternativess_(1)s_(2)
Option 1$10,000$30,000
Option 2$5,000$45,000
Option 3$-4,000$60,000


A. $60,000; Option 3 

B. $28,000, Option 3 

C. $25,000: Option 2 

D. $20,000 Option 1 

E. $32,000: Option 3
Answer #1

Equally likely criterion views each state as equal and weighs them equally. Here we have two states so we assign (0.5) weight to each and add both the states. Then choose the state with maximum value.

Option 1 = 10000(0.5) + 30000(0.5) =5000+15000 = $20,000

Option 2 = 5000(0.5) + 45000(0.5) =2500+22500 = $25,000

Option 3 = - 4000(0.5) + 60000(0.5) = -2000 + 30000 = $28,000

Hence, the highest value for equally likely is $28,000 for option 3.

The correct option is B.

Do give a positive rating if you think this answer helped